Determine the mass of nitrogen found in 8.85g of Pb(NO3)3


1
Expert's answer
2022-01-25T10:55:03-0500

Solution:

Pb(NO3)3 is not the correct chemical formula.

Pb(NO3)2 is the correct chemical formula.

Considering the above, write down the chemical scheme:

Pb(NO3)2 → Pb + 2N + 6N

According to the scheme:

Moles of Pb(NO3)2 = Moles of N / 2


The molar mass of Pb(NO3)2 is 331.2 g/mol

Hence,

Moles of Pb(NO3)2 = (8.85 g) × (1 mol / 331.2 g) = 0.02672 mol


Moles of N = 2 × Moles of Pb(NO3)2 = 2 × 0.0267 mol = 0.05344 mol

The atomic weight of N is 14.0067 g/mol

Hence,

Mass of N = (0.05344 mol) × (14.0067 g / 1 mol) = 0.7485 g = 0.75 g

Mass of N = 0.75 g


Answer: The mass of nitrogen (N) is 0.75 grams